# Backdoor defense, learnability and obfuscation

### v1

@@ 1 Introduction
Our formal notion is focused on backdoor detection at runtime, meaning that the defender is given a particular input and must flag whether or not it activates the backdoor trigger. We focus on this case because of the existence of undetectable backdoors if the defender is instead required to flag a model as backdoored without being given any particular input (Goldwasser et al., 2022). Moreover, detection at runtime is sufficient in threat models where the attacker is given only one opportunity to modify the model, being akin to giving the defender an additional chance to modify the model after the attacker. We also focus on the white-box setting, in which the defender has access to a complete description of the model.
@@ 2 Related work
The most closely related work to our own is that of Goldwasser et al. (2022), whose main results use a digital signature scheme to insert an undetectable backdoor into a model. The backdoor trigger can be applied to any input by perturbing it appropriately. In the black-box setting, the backdoor is undetectable in the sense that it is computationally infeasible for the defender to find any input on which the backdoored model and the original model differ. In the white-box setting, the backdoor is undetectable in the sense that it is computationally infeasible for the defender to distinguish the backdoored model from the original model. However, both of these results are in the setting where the backdoor trigger must be extracted from the model rather than detected at runtime.
In the defense at runtime setting, Goldwasser et al. (2022) also show that a backdoored model can be modified by the defender to produce a model that is not backdoored. Once again, the backdoor trigger is applied by perturbing the input. The result only applies to models that satisfy a certain smoothness assumption, and mirrors the use of randomized smoothing to achieve certified adversarial robustness by Cohen et al. (2019). By contrast, we avoid making smoothness assumptions, and further develop the theory of runtime backdoor detection.
@@
Our result on statistical defendability is heavily based on the result of Hanneke et al. (2022) in the realizable case. Following Levine and Feizi (2020) and Jia et al. (2021), they use a voting algorithm to produce a backdoor defense in the data-poisoning setting. By combining this with a classical result of Haussler et al. (1994b) on optimal learning, they relate the optimal performance of this algorithm to the VC dimension of the function class. Our result adapts this to our more general notion of defendability.
@@ 4 Statistical defendability
In their classical paper on prediction strategies, Haussler et al. (1994b) exhibit a “1-inclusion graph” algorithm yielding the following result.
@@ 5.3 Efficient defendability and obfuscation
Here, $\mathsf{OWF}$ again denotes the existence of a one-way function, and $\mathsf{iO}$ denotes the existence of an efficient indistinguishability obfuscator. Roughly speaking, an efficient indistinguishability obfuscator is a probabilistic polynomial-time algorithm that takes in a circuit and outputs an “obfuscated” circuit with the same behavior. The circuit being “obfuscated” means that the obfuscations of two different circuits with the same behavior are not distinguishable by any probabilistic polynomial-time adversary. For a precise definition of an efficient indistinguishability obfuscator, we refer the reader to Sahai and Waters (2014, Section 3). The study indistinguishability obfuscation was initiated by Barak et al. (2001), and an efficient indistinguishability obfuscator was constructed from well-studied computational hardness assumptions by Jain et al. (2021).
@@ 6 Defendability of decision trees
In our study of decision trees, we focus on the uniform-PAC model of learning, in which the distribution $\mathcal{D}$ is always the uniform distribution. Thus we say that a representation class is efficiently uniform-PAC learnable to mean that it is efficiently PAC learnable as long as $\mathcal{D}$ is uniform. Similarly, we say that a representation class is efficiently uniform-defendable to mean that it is efficiently defendable as long as $\mathcal{D}$ is uniform. Note that efficient uniform PAC-learnability implies efficient uniform defendability, by specializing the proof of Corollary 5.1.
As far as we are aware, it is unknown whether the representation class of polynomial size decision trees over $\left\{0,1\right\}^{n}$ is efficiently uniform-PAC learnable. Polynomial-time PAC learning algorithms are known for certain product distributions (Kalai and Teng, 2008), as well as for the uniform distribution if we allow membership queries (i.e., calls to an oracle that outputs the value of the function on an input of the algorithm’s choice) (Goldreich and Levin, 1989; O’Donnell, 2021, Chapter 3.5). Such algorithms generally use at least linearly many calls to the example or membership oracle.
By contrast, we now show that this representation class is efficiently uniform-defendable using a very fast detection strategy.
@@ A.2 Conditions for accuracy
In this section, we give bounds on the error of the posterior. Our results are closely related to the results of Haussler et al. [1994a]; but they give bounds on the error that an optimal learning algorithm gets after seeing $m$ examples, whereas we give bounds on the error that a Boltzmann posterior gets with an inverse temperature of $m$. These models come apart in the sense that in their setting, the optimal error rate for classes of VC-dimension $d$ is $\Theta\mathopen{}\left(\frac{m}{d}\right)$ (due to a result in Haussler et al. [1994b]), whereas in our setting the optimal error rate is $\Theta\mathopen{}\left(\frac{m}{d}\log\frac{d}{m}\right)$. On an intuitive level, this is because the Boltzmann posterior never “fully updates”: it is more akin to the posterior one might get after drawing $m$ examples that have a constant amount of classification noise, leaving some uncertainty.
@@ A.2.3 Matching lower bound
The following result shows that the bound $O\mathopen{}\left(\frac{d}{m}\log\frac{m}{d}\right)$ we obtained in Lemmas A.5 and A.4 is tight, which means that the best bound on $\delta$ that can be obtained by any approach based on Lemma A.1 is $\Theta\mathopen{}\left(d\varepsilon\log\frac{1}{d\varepsilon}\right)$.

### v2

@@ 1 Introduction
Our formal notion is focused on backdoor detection at runtime, meaning that the defender is given a particular input and must flag whether or not it activates the backdoor trigger. We focus on this case because of the existence of undetectable backdoors if the defender is instead required to flag a model as backdoored without being given any particular input (Goldwasser, Kim, Vaikuntanathan, and Zamir, 2022). Moreover, detection at runtime is sufficient in threat models where the attacker is given only one opportunity to modify the model, being akin to giving the defender an additional chance to modify the model after the attacker. We also focus on the white-box setting, in which the defender has access to a complete description of the model.
@@ 2 Related work
The most closely related work to our own is that of Goldwasser, Kim, Vaikuntanathan, and Zamir (2022), whose main results use a digital signature scheme to insert an undetectable backdoor into a model. The backdoor trigger can be applied to any input by perturbing it appropriately. In the black-box setting, the backdoor is undetectable in the sense that it is computationally infeasible for the defender to find any input on which the backdoored model and the original model differ. In the white-box setting, the backdoor is undetectable in the sense that it is computationally infeasible for the defender to distinguish the backdoored model from the original model. However, both of these results are in the setting where the backdoor trigger must be extracted from the model rather than detected at runtime.
In the defense at runtime setting, the same authors also show that a backdoored model can be modified by the defender to produce a model that is not backdoored. Once again, the backdoor trigger is applied by perturbing the input. The result only applies to models that satisfy a certain smoothness assumption, and mirrors the use of randomized smoothing to achieve certified adversarial robustness by Cohen et al. (2019). By contrast, we avoid making smoothness assumptions, and further develop the theory of runtime backdoor detection.
Concurrent work by Goldwasser, Shafer, Vafa, and Vaikuntanathan (2024) studies a black-box, non-realizable version of backdoor defense at runtime. Rather than predicting the exact label of the non-backdoored model on the given input (which is equivalent to our notion of runtime detection), the defender must instead produce a “canonical” label that has high accuracy on average, by using black-box access to the potentially backdoored model. In this local mitigation setting, they construct efficient defenders for almost-linear and almost-polynomial functions. They also consider the global mitigation setting, in which the defender must produce an entire clean model, constructing efficient global mitigators for Fourier-heavy functions.
@@
Our result on statistical defendability is heavily based on the result of Hanneke et al. (2022) in the realizable case. Following Levine and Feizi (2020) and Jia et al. (2021), they use a voting algorithm to produce a backdoor defense in the data-poisoning setting. By combining this with a classical result of Haussler, Littlestone, and Warmuth (1994b) on optimal learning, they relate the optimal performance of this algorithm to the VC dimension of the function class. Our result adapts this to our more general notion of defendability.
@@ 4 Statistical defendability
In their classical paper on prediction strategies, Haussler, Littlestone, and Warmuth (1994b) exhibit a “1-inclusion graph” algorithm yielding the following result.
@@ 5.3 Efficient defendability and obfuscation
Here, $\mathsf{OWF}$ again denotes the existence of a one-way function, and $\mathsf{iO}$ denotes the existence of an efficient indistinguishability obfuscator. Roughly speaking, an efficient indistinguishability obfuscator is a probabilistic polynomial-time algorithm that takes in a circuit and outputs an “obfuscated” circuit with the same behavior. The circuit being “obfuscated” means that the obfuscations of two different circuits with the same behavior are not distinguishable by any probabilistic polynomial-time adversary. For a precise definition of an efficient indistinguishability obfuscator, we refer the reader to Sahai and Waters (2014, Section 3). The study indistinguishability obfuscation was initiated by Barak et al. (2001), and an efficient indistinguishability obfuscator was constructed from well-studied computational hardness assumptions by Jain, Lin, and Sahai (2021).
@@ 6 Defendability of decision trees
In our study of decision trees, we focus on the uniform-PAC model of learning, in which the distribution $\mathcal{D}$ is always the uniform distribution. Thus we say that a representation class is efficiently uniform-PAC learnable to mean that it is efficiently PAC learnable as long as $\mathcal{D}$ is uniform. Similarly, we say that a representation class is efficiently uniform-defendable to mean that it is efficiently defendable as long as $\mathcal{D}$ is uniform.
As far as we are aware, it is unknown whether the representation class of polynomial size decision trees over $\left\{0,1\right\}^{n}$ is efficiently uniform-PAC learnable, although polynomial-time PAC learning algorithms are known for certain product distributions (Kalai and Teng, 2008). However, the class is efficiently uniform PAC-learnable if we allow membership queries (i.e., calls to an oracle that outputs the value of the function on an input of the algorithm’s choice) (Goldreich and Levin, 1989; O’Donnell, 2021, Chapter 3.5). By specializing the proof of Corollary 5.2, it follows that the representation class of polynomial size decision trees over $\left\{0,1\right\}^{n}$ is efficiently defendable, since the defender can efficiently compute the results of membership queries for themselves using the representation of the decision tree.
Nevertheless, these learning algorithms for decision trees generally use at least linearly many calls to the example or membership oracle. By contrast, we now show that there is a defense for decision trees that is much faster than this.
@@ A.2 Conditions for accuracy
In this section, we give bounds on the error of the posterior. Our results are closely related to the results of Haussler, Kearns, and Schapire [1994a]; but they give bounds on the error that an optimal learning algorithm gets after seeing $m$ examples, whereas we give bounds on the error that a Boltzmann posterior gets with an inverse temperature of $m$. These models come apart in the sense that in their setting, the optimal error rate for classes of VC-dimension $d$ is $\Theta\mathopen{}\left(\frac{m}{d}\right)$ (due to a result in Haussler, Littlestone, and Warmuth [1994b]), whereas in our setting the optimal error rate is $\Theta\mathopen{}\left(\frac{m}{d}\log\frac{d}{m}\right)$. On an intuitive level, this is because the Boltzmann posterior never “fully updates”: it is more akin to the posterior one might get after drawing $m$ examples that have a constant amount of classification noise, leaving some uncertainty.
@@ A.2.3 Matching lower bound
The following result shows that the bound $O\mathopen{}\left(\frac{d}{m}\log\frac{m}{d}\right)$ we obtained in Lemmas A.15 and A.11 is tight, which means that the best bound on $\delta$ that can be obtained by any approach based on Lemma A.6 is $\Theta\mathopen{}\left(d\varepsilon\log\frac{1}{d\varepsilon}\right)$.